### v1

# Callan-Symanzik-like equation in information theory

Abstract
In this paper, we consider the jumps in the complexity growth rate (CGR) through the complexity=anything proposal. It is revealed that bulk fields are in charge of the height and the location of these jumps. To put it accurately, their counterparts are divergence of energy momentum tensor, shear viscosity and the Weyl anomaly of the boundary theory. The behavior of CGR around these jumps manifests a critical behavior and critical exponents. Moreover, it is figured out that CGR satisfies a Callan-Symanzik-like equation around these jumps. This equation suggests that the speed of information processing could be changed by the scale of energy. Put it another way, by change of the scale of energy one could turn the classical computers into faster processor computers even faster than our current quantum computers.
1 Introduction
The AdS/CFT duality establishes a profound correspondence between a gravitational theory in the bulk of anti-de Sitter (AdS) space and a conformal field theory (CFT) on its boundary [1]. This holographic principle provides a powerful dictionary, allowing quantities that are computationally intractable in one theory—particularly in strongly coupled regimes—to be studied through their dual counterparts in the other. A striking example arises in the context of black hole physics, where the growth of Einstein-Rosen bridge volume in the bulk poses a conceptual puzzle: if a black hole is in thermodynamic equilibrium, and its dual boundary theory is likewise equilibrated, what physical meaning should be ascribed to this increasing volume? Recent insights suggest that this geometric evolution corresponds to the computational complexity of the boundary theory, specifically the minimal number of fundamental operations required to prepare a target quantum state from a given reference state [2].
In holography, several conjectures relate bulk geometric quantities to boundary complexity. Among these proposals three of them have been widely explored, the complexity=volume (CV) in which the volume of Einstein-Rosen bridges are corresponded to the complexity of the boundary theory, [3], complexity=action (CA), relates the action of the bulk theory restricted to the Wheeler-DeWitt patch to the complexity [4] and complexity=volume (CV2.0) corresponds the spacetime volume of the bulk restricted to the Wheeler-DeWitt patch to the complexity [5]. More recently, a generalized framework—dubbed complexity=anything—has emerged, positing that arbitrary bulk volumes can encode boundary complexity, offering a broader and more flexible perspective [6]. Intriguingly, the complexity growth rate (CGR) under this proposal exhibits distinct branches, which some interpretations identify them as signatures of phase transitions [7, 8].
This work investigates the physical origins of these jumps, addressing two key questions: (1) What determines their heights? (2) What controls their temporal locations? We demonstrate that the jumps positions are dynamically influenced by bulk fields, which, via the fluid-gravity duality, map to specific boundary quantities: the energy-momentum tensor, Weyl anomaly [9] and shear viscosity [10]. Consequently, the location of these transitions depends on the divergence of the boundary fluid’s energy-momentum tensor (indicating energy injection), its shear viscosity, and its anomalous conformal behavior. Meanwhile, the jump amplitudes are linked to the renormalization group (RG) flow of the boundary theory.
Remarkably, the CGR transitions exhibit critical phenomena, characterized by scaling behavior and universality. Near these critical points, the system obeys a Callan-Symanzik-like equation, where the beta function is expressed in terms of the generalization parameter from the complexity=anything proposal, alongside an anomalous dimension. This criticality further displays universal features, independent of the specific choice of the generalized parameter.
The paper is structured as follows: the section 2 introduces the complexity=anything framework and defines CGR for a generic bulk metric. the section 3 elucidates the connection between jump locations and bulk field dynamics. The section 4 analyzes the critical behavior of CGR, derives the associated critical exponents, and presents the Callan-Symanzik-like equation in the context of information theory. We conclude with a discussion of implications in quantum computers and future directions.
2 Different branches in $C_{Any}$
Recently, $\mathcal{C}_{Any}$ (complexity equals to anything) conjecture has emerged as a generalized version of complexity in the bulk. In this proposal the complexity is not limited to a specific geometric quantity but can be described by a more general function to include a class of new diffeomorphism invariant observables [6]:
$\mathcal{O}_{F_{1},\Sigma_{F_{2}}}(\Sigma_{CFT})=\frac{1}{G_{N}L}\int_{\Sigma_{F_{2}}}\mathrm{d}^{d}\sigma\sqrt{h}F_{1}(g_{\mu\nu};X^{\mu})$ (1)
where, $F_{1}$ can be a general scalar function of metric $g_{\mu\nu}$ and an embedding $X^{\mu}(\sigma^{a})$ of the hypersurfaces. Furthermore, $\Sigma_{F_{2}}$ is codimension-one hypersurface in the bulk spacetime with boundary time slice $\partial\Sigma_{F_{2}}=\Sigma_{CFT}$. Extremality of the hypersurface leads to:
$\delta_{X}\Big{(}\int_{\Sigma}\mathrm{d}^{d}\sigma\sqrt{h}F_{2}(g_{\mu\nu};X^{\mu})\Big{)}=0.$ (2)
For simplicity, we follow the case $F_{1}=F_{2}$, so the observable (1), known by $\mathcal{C}_{Any}$, obeying above condition is expressed by:
$\mathcal{C}_{Any}(\tau)=\max_{\partial\Sigma(\tau)=\Sigma_{\tau}}\frac{V_{x}}{G_{N}L}\left[\int_{\Sigma}\mathrm{d}^{d}\sigma\sqrt{h}F_{1}(g_{\mu\nu};X^{\mu})\right],$ (3)
where $h$ is the determinant of induced metric on the given hypersurface. One of the usual choice of generalization is:
$F_{1}=F_{2}=1+\gamma L^{4}\mathcal{C}^{2},$ (4)
where $\mathcal{C}^{2}$ is the Weyl tensor squared and for $\gamma=0$ the CV is recovered. In the following, we are going to study the $\mathcal{C}_{Any}$ of a generic metric:
$\mathrm{d}s^{2}=g_{tt}\mathrm{d}t^{2}+g_{rr}\mathrm{~d}r^{2}+\frac{r^{2}}{L^{2}}d\Omega^{2}.$ (5)
To done this, we need to transform the metric to Eddington-Finkelstein coordinates:
$\mathrm{d}s^{2}=g_{tt}\mathrm{d}v^{2}-2\sqrt{-g_{tt}g_{rr}}\mathrm{~d}v\mathrm{~d}r+\frac{r^{2}}{L^{2}}d\Omega^{2},$ (6)
where:
$v=t+r^{*}(r),\quad\mathrm{d}r^{*}=\sqrt{-\frac{g_{rr}}{g_{tt}}}\mathrm{d}r.$ (7)
Then, the complexity could be computed from (3) and (4) as [6]:
$\mathcal{C}_{Any}(\tau)=\frac{V_{0}}{G_{N}L}\int_{\Sigma}\mathrm{d}\sigma r^{d-1}\sqrt{g_{tt}\dot{v}^{2}-2\sqrt{-g_{tt}g_{rr}}\dot{v}\dot{r}}a(r),$ (8)
where the dots indicate derivatives with respect to $\sigma$, and $V_{0}$ represents the volume of spatial directions.
Supposing $\mathcal{C}_{Any}$ as an action leads to a conserved momentum conjugate to coordinate $v$, because of spacetime is stationary:
$P_{v}=-\frac{\partial\mathcal{L}}{\partial\dot{v}}=-\frac{a(r)r^{d-1}(g_{tt}\dot{v}-\sqrt{-g_{tt}g_{rr}}\dot{r})}{\sqrt{g_{tt}\dot{v}^{2}-2\sqrt{-g_{tt}g_{rr}}\dot{v}\dot{r}}}=\frac{r^{d-1}}{H(r)}\Big{(}\dot{r}-\sqrt{-\frac{g_{tt}}{g_{rr}}}\dot{v}\Big{)}.$ (9)
Due to the fact that (8) is diffeomorphism invariant and doesn’t change under reparametrization, the parameter $\sigma$ could be fixed by choosing:
$\sqrt{g_{tt}\dot{v}^{2}-2\sqrt{-g_{tt}g_{rr}}\dot{v}\dot{r}}=a(r)\sqrt{-g_{tt}g_{rr}}H(r).$ (10)
Then it is straightforward to derive extremality conditions:
$\dot{r}=\pm\sqrt{\frac{H(r)^{2}P^{2}_{v}}{r^{2(d-1)}}-g_{tt}a(r)^{2}H(r)^{2}},$ (11)
$\dot{t}=\dot{v}-\sqrt{-\frac{g_{rr}}{g_{tt}}}\dot{r}=\frac{-\sqrt{-\frac{g_{rr}}{g_{tt}}}P_{v}\dot{r}H(r)}{r^{d-1}\sqrt{\frac{H(r)^{2}P^{2}_{v}}{r^{2(d-1)}}-U(r)}},$ (12)
where $U(r)$ is an effective potential.
$U(r)=g_{tt}a(r)^{2}H(r)^{2}.$ (13)
Then, (11) gives the equation of motion like for a classical particle:
$\dot{r}^{2}+U(r)=\frac{H(r)^{2}P^{2}_{v}}{r^{2(d-1)}}.$ (14)
In the symmetric trajectory, the conserved momentum is a function of turning point $r_{min}$, we could redefine the potential as follows:
$\tilde{U}(r)=g_{tt}a(r)^{2}r^{2(d-1)}.$ (15)
Then:
$P^{2}_{v}=\tilde{U}(r_{min})=g_{tt}(r_{min})a(r_{min})^{2}r_{min}^{2(d-1)}.$ (16)
Boundary time derivation of $\mathcal{C}_{Any}$ leads to:
$\frac{\mathrm{d}\mathcal{C}_{Any}}{\mathrm{~d}\tau}=\frac{1}{2}\frac{\mathrm{d}\mathcal{C}_{Any}}{\mathrm{~d}\tau_{\mathrm{R}}}$ $=\frac{V_{0}}{G_{N}L}P_{v}$
$=\frac{V_{0}}{G_{N}L}\sqrt{g_{tt}(r_{min})}a(r_{min})r_{min}^{(d-1)}.$ (17)
Then, the growth rate of the generalized complexity can be studied by the behavior of the conserved momentum in full-time. The late-time behavior of the growth rate is determined by:
$\lim_{\tau\rightarrow\infty}\frac{\mathrm{d}\mathcal{C}_{Any}}{\mathrm{~d}\tau}=\frac{V_{0}}{G_{N}L}\sqrt{g_{tt}(\tilde{r}_{\min})}a(\tilde{r}_{\min})\tilde{r}_{min}^{(d-1)},$ (18)
where $\tilde{r}_{\min}$ is local maximum of the effective potential, and because at this radius time goes to infinity, it is known sometimes as $r_{f}:=\tilde{r}_{\min}$.
3 What are in charge of these jumps in CGR?
It has been seen that the complexity equals anything shows some jumps in the complexity growth rate while the former proposals do not. This phenomenon raises the question that why there are these jumps in this proposal and what they refer to. Are they related to a physical phenomena? In this section we answer to the questions affirmatively and seeking the reason on the boundary theory of the bulk.
The jump occurs around the asymptotic behavior of CGR, the extrimum values of $P_{v}$, when the potential (15) has a local maxima, in other words $\frac{d\tilde{U}}{dr}=0$. For a generic solution we have got:
$ds^{2}=$ $g_{tt}dt^{2}+g_{rr}dr^{2}+r^{2}d\Omega^{2}$ (19)
$\tilde{U}$= $g_{tt}a^{2}r^{2(d-1)}$ (20)
$g_{tt}a^{2}r^{2(d-1)}=cte:=\alpha.$ (21)
In the following we set $\alpha=1$. By rescaling $g_{tt}\rightarrow r^{-2(d-1)-2}g_{tt}=g_{t^{\prime}t^{\prime}}$ and $dt^{2}\rightarrow r^{2(d-1)+2}dt^{2}=dt^{\prime 2}$ one could rewrite (21) as following:
$r^{-2}g_{tt}a^{2}$ $=1$ (22)
$r^{-2}g_{tt}(1+\gamma C^{2})^{2}$ $=1.$ (23)
At the leading term in terms of $\gamma$ [^1] We suppose that $\gamma\ll 1$.:
$r^{-2}g_{tt}+2\gamma r^{-2}g_{tt}C^{2}$ $=1$ (24)
If one writes the bulk Weyl squared tensor in terms of the boundary coordinates:
$\mathcal{C}_{\mu\nu\rho\sigma}\mathcal{C}^{\mu\nu\rho\sigma}=\mathcal{C}_{ijkl}\mathcal{C}^{ijkl}+4\mathcal{C}_{i\rho j\rho}\mathcal{C}^{i\rho j\rho}+4\mathcal{C}_{ijk\rho}\mathcal{C}^{ijk\rho},$ (25)
where ijkijk... are boundary coordinates and $\rho$ the radial. Using the Fefferman-Graham coordinate to write the bulk Weyl squared tensor in terms of the boundary values it is given by (see Appendix A)  [^2] It should be noted that we used the FG coordinates at the finite cut-off $r=r_{j}$ where $r_{j}$ is a location of a local maxima of the potential.:
$C^{2}$ $=\mathscr{C}+\Big{(}<T_{ij}>-\frac{g_{ij}^{(0)}}{4}<T_{k}^{k}>\Big{)}^{2}+\Big{(}\nabla_{i}<T_{jk}>\Big{)}^{2}+\mathcal{O}(\rho^{2}),$ (26)
$=C_{(0)}^{2}+\mathcal{O}(\rho^{2}),$ (27)
where $\mathscr{C}^{2}$ is the boundary Weyl tensor squared, $T_{ij}>$ the expectation value of the boundary energy-momentum tensor and $g_{ij}^{(0)}$ the induced metric on the boundary. The first term in the right-hand-side of (27) is the Weyl anomaly of the boundary theory [9], the second term is related to the shear viscosity of the boundary theory [10] and the third term is not conserved in the presence of matter that shows the flow of the matter through the boundary [11]. By rewriting (24) in the FG coordinates one gets [^3] The coordinate change in (23) makes the following formula, (28), to be written in terms of the first order of the Weyl squared tensor expansion. If one does not use the coordinate change, (28) is written in terms of the other orders of the Weyl squared tensor expansion. However, we know that the other terms could be rewritten in terms of the first order. Then, the coordinate change makes the result more clear. In addition, we should note that $r_{i}$ appears due to the FG coordinates at $r_{i}$ cut-off:
$g^{(0)}_{tt}+2\gamma g^{(0)}_{tt}\Bigg{(}\mathscr{C}+\Big{(}<T_{ij}>-\frac{g_{ij}^{(0)}}{4}<T_{k}^{k}>\Big{)}^{2}+\Big{(}\nabla_{i}<T_{jk}>\Big{)}^{2}+\mathcal{O}(\rho^{2})\Bigg{)}=1.$ (28)
(28) shows that the jumps in the complexity are refereed to the following quantities in the boundary theory:
  • The Weyl anomaly, $\mathscr{C}$.
  • The shear viscosity, $\Big{(}<T_{ij}>-\frac{g_{ij}^{(0)}}{4}<T_{k}^{k}>\Big{)}^{2}$.
  • Non-conservation of the energy-momentum on the boundary, $\Big{(}\nabla_{i}<T_{jk}>\Big{)}^{2}$.
  • The boundary metric, $g^{(0)}_{tt}$.
  • The parameter of the generalization, $\gamma$.
If there is no Weyl anomaly in the boundary theory and no shear viscosity, then it is expected that the jumps in the complexity are related to:
$C_{(0)}^{2}=\Big{(}\nabla_{i}<T_{jk}>\Big{)}^{2}=\frac{1-g^{(0)}_{tt}}{2\gamma g^{(0)}_{tt}}$. (29)
In [11] it has been shown that the non-conservation of the energy-momentum tensor on the boundary is related to the presence of the derivative of a bulk field, say a scalar field:
$\nabla^{i}<T_{ij}>=<\hat{O}>\partial_{j}\phi^{(0)},$ (30)
where $\mathcal{O}$ is the operator sourced by the bulk field $\phi$. It has been revealed that the value of the bulk fields on the boundary could determine the location of the jumps in CGR. Having done that, there would be a question that if the distance of the jumps is related to a physical quantity or not? The distance between the jumps could be realized by $d_{jumps}=|r_{i}-r_{j}|$ where $r_{i}$ means the location where the ith"i"th jump occurs at. As it has been studied, the location of the jump is where the potential $\tilde{U}$ has a local maxima. In other words, roots of (21). The distance between roots depends on the quantities in (21), quantities in $g_{tt}$ and $\gamma$ the coefficient of the generalization:
$g_{tt}a^{2}r^{-2}=1$ (31)
$Q(r_{i})=g_{tt}a^{2}r^{-2}-1$ (32)
where $r_{i}$ is the ith"i"th root of (21). The dependence could be regarded as:
$\frac{\partial r_{i}}{\partial a}=\frac{\frac{\partial Q}{\partial a}}{\frac{\partial Q}{\partial r}}\Big{|}_{r=r_{i}}=\frac{\partial Q}{\partial a}\Big{|}_{r=r_{i}}=\frac{\partial\tilde{U}}{\partial a}\Big{|}_{r=r_{i}}$ (33)
where $\frac{\partial r_{i}}{\partial a}$ determines the dependency of the distance to $a$, one the quantities in $g_{tt}$ and $\gamma$. As an example we show that how jumps are controlled by the parameter of the generalization:
$\frac{\partial r_{i}}{\partial\gamma}=2\Big{(}g_{tt}^{(0)}C_{(0)}^{2}\Big{)}\Big{|}_{r=r_{i}}.$ (34)
As a consequence the distance between jumps could be written:
$\frac{\partial\big{(}r_{i}-r_{j}\big{)}}{\partial\gamma}=2\Bigg{(}g_{tt}^{(0)}C_{(0)}^{2}\Big{|}_{r=r_{i}}-g_{tt}^{(0)}C_{(0)}^{2}\Big{|}_{r=r_{j}}\Bigg{)},$ (35)
which means that the distance between jumps is encoded in the RG flow of the boundary theory i.e. the divergence of the energy-momentum tensor, shear viscosity and the Weyl anomaly at different energy scales. Moreover, (35) tells about the scale-dependent origin of the distance of the jumps in CGR.
4 Critical exponents
As it has been revealed that the location of these jumps are scale dependent. This result suggests that the CGR itself could behave differently in various scales. In this section we are going to study the behavior of CGR in different scales. In this manner CGR is a function of the generalization parameter; however, the location of these jumps $r_{i}$ is a function of the parameter, so we could use a chain rule:
$\frac{d\dot{\mathcal{C}}}{d\gamma}\Big{|}_{r_{i}}=$ $\frac{\partial\dot{\mathcal{C}}}{\partial\gamma}\Big{|}_{r_{i}}+\frac{\partial\dot{\mathcal{C}}}{\partial r}\frac{\partial r}{\partial\gamma}\Big{|}_{r_{i}}$ (36)
$=$ $\frac{\partial\dot{\mathcal{C}}}{\partial\gamma}\Big{|}_{r_{i}}+2g_{tt}^{(0)}C_{(0)}^{2}\frac{\partial\dot{\mathcal{C}}}{\partial r}\Big{|}_{r_{i}}.$ (37)
The above result appears when we are working in $t^{\prime}$ coordinate and Fefferman-Graham coordinate at $r_{i}$ cut-off. One could proceed with $t$ coordinate and $r_{i}$ cut-off, then:
$\frac{d\dot{\mathcal{C}}}{d\gamma}\Big{|}_{r_{i}}=\frac{\partial\dot{\mathcal{C}}}{\partial\gamma}\Big{|}_{r_{i}}+2r_{i}^{2(d)}g_{tt}^{(0)}C_{(0)}^{2}\frac{\partial\dot{\mathcal{C}}}{\partial r}\Big{|}_{r_{i}}.$ (38)
Around $r_{i}$, where it is a late time and asymptotic behavior of $\dot{\mathcal{C}}$ in Fefferman-Graham coordinates at $r_{i}$ cut-off, CGR at the leading order could be written as:
$\dot{\mathcal{C}}\sim P_{v}=\sqrt{\tilde{U}}$ $=\sqrt{g_{tt}a^{2}r^{2(d-1)}}$ (39)
$=\Bigg{(}r_{i}^{2(d-1)}\Big{(}r_{i}^{2}g_{tt}^{(0)}+g_{tt}^{(1)}+\frac{g_{tt}^{(2)}}{r_{i}^{2}}+...\Big{)}\Big{(}C_{(0)}^{2}+\frac{C_{(1)}^{2}}{r_{i}^{2}}+...\Big{)}\Bigg{)}^{\frac{1}{2}}$ (40)
$\approx\Big{(}r_{i}^{2d}g_{tt}^{(0)}+r_{i}^{2d}(2\gamma g_{tt}^{(0)}C_{(0)}^{2})\Big{)}^{\frac{1}{2}}.$ (41)
As a consequence, if $\gamma\rightarrow\lambda^{\nu}\gamma$ and $r_{i}\rightarrow\lambda^{\mu}r_{i}$, then $\dot{\mathcal{C}}(\lambda^{\mu}r_{i},\lambda^{\nu}\gamma)=\lambda^{\Delta}\dot{\mathcal{C}}(r_{i},\gamma)$, where $\Delta$, $\mu$ and $\nu$ are critical exponents. Derivative of CGR with respect to $\lambda$ and setting $\lambda=1$ is given:
$\frac{d\dot{\mathcal{C}}}{d\lambda}\Big{|}_{\lambda=1}$ $=\frac{\partial\dot{\mathcal{C}}}{\partial r_{i}}\frac{\partial r_{i}}{\partial\lambda}+\frac{\partial\dot{\mathcal{C}}}{\partial\gamma}\frac{\partial\gamma}{\partial\lambda}$ (42)
$=\frac{\partial\dot{\mathcal{C}}}{\partial r_{i}}(\mu r_{i})+\frac{\partial\dot{\mathcal{C}}}{\partial\gamma}(\nu\gamma).$ (43)
In addition:
$\frac{d\dot{\mathcal{C}}(\lambda^{\mu}r_{i},\lambda^{\nu}\gamma)}{d\lambda}\Big{|}_{\lambda=1}=\frac{d}{d\lambda}\Big{(}\lambda^{\Delta}\dot{\mathcal{C}}(r_{i},\gamma)\Big{)}\Big{|}_{\lambda=1}=\Delta\dot{\mathcal{C}}(r_{i},\gamma).$ (44)
Combining (43) and (44):
$\Big{(}(2r_{i}^{2d}\nu\gamma g_{tt}^{(0)}C_{(0)}^{2}-\mu r_{i})\frac{\partial}{\partial r_{i}}-\nu\gamma\frac{d}{d\gamma}+\Delta\Big{)}\dot{\mathcal{C}}=0,$ (45)
where $(2r^{2d}\nu\gamma g_{tt}^{(0)}C_{(0)}^{2}-\mu r_{i})\frac{\partial\dot{\mathcal{C}}}{\partial r}$ shows how CGR is changed by the scale of the energy $r_{i}$, $\nu\gamma\frac{d\dot{\mathcal{C}}}{d\gamma}$ indicating how CGR is changed by the generalization parameter, in other words, if the CGR is increasing by increase in $\gamma$ or decreasing. We call (55) Callan-Symanzik-like equation in information theory, where $r_{i}$ is the information renormalization scale, $\beta=-\nu\gamma$ beta function in information theory and $\Delta$ information anomalous dimension.
(41) suggests that CGR near $r_{i}$ is scaled:
$\dot{\mathcal{C}}(\lambda^{\mu}r_{i},\lambda^{\nu}\gamma)$ $\sim\Big{(}\lambda^{2d\mu}r_{i}^{2d}g_{tt}^{(0)}+\lambda^{2d\mu+\nu}r_{i}^{2d}(2\gamma g_{tt}^{(0)}C_{(0)}^{2})\Big{)}^{\frac{1}{2}}$ (46)
$=\lambda^{\Delta}\dot{\mathcal{C}}(r_{i},\gamma)$ (47)
Now we could consider the validity of critical exponents:
$\Delta=d\mu$ (48)
$\Delta=d\mu+\frac{\nu}{2}$ (49)
The solution for the critical exponents would be $\nu=0$ and $\Delta=d\mu$. This solution demonstrates a universal critical behavior around the $r_{i}$’s or jumps of CGR.
However, for large $\gamma$ we could write $a^{2}\approx 1+\gamma^{2}C^{4}$. Then (41) is rewritten as:
$\dot{\mathcal{C}}\sim P_{v}=\sqrt{g_{tt}a^{2}r^{2(d-1)}}=\Big{(}r_{i}^{2d}g_{tt}^{(0)}+r_{i}^{2d}(\gamma^{2}g_{tt}^{(0)}(C_{(0)}^{2})^{2})\Big{)}^{\frac{1}{2}}.$ (51)
The critical exponents would be:
$\Delta=d\mu$ (52)
$\Delta=d\mu+\nu$ (53)
For the large $\gamma$ the same solution appears. This result illustrates a universal critical behavior around the phase transition of CGR regardless of the generalization parameter $\gamma$. Looking again at (55) by the solution reveals:
$\Big{(}-\mu r_{i}\frac{\partial}{\partial r_{i}}+d\mu\Big{)}\dot{\mathcal{C}}=0,$ (55)
The solution of the above equation would be:
$\dot{\mathcal{C}}\sim r_{i}^{d}(\gamma)+\alpha,$ (56)
where $\alpha$ is a constant in (21). By the fact that $\dot{\mathcal{C}}$ is constant at $r_{i}$, then it could be written as:
$\dot{\mathcal{C}}\sim\Big{(}r-r_{i}(\gamma)\Big{)}^{d}+\alpha.$ (57)
This solution shows that the critical behavior of CGR around the critical point $r_{i}$. Then, the critical exponents $\Delta=d\mu=d$.
5 Conclusion
The AdS/CFT by corresponding a gravitational theory to a quantum theory could alleviate the perturbation calculation in strongly-coupled systems. Exploring the counterpart quantities in the AdS/CFT make it a strong technique to study the physical theories. Computational complexity as one of these quantities has been considered in holography and provided different proposal. Recently, a proposal known as complexity equals anything has been introduced in which the volume of a generalized surface is corresponded to the complexity of the boundary theory. Studying the complexity of various models through this proposal shows some branches in CGR. Moreover, these branches has been considered as a phase transition in CGR.
Because of the fact that the phase transition appeared in complexity equals anything, this suggests that the generalized surfaces could make the transition. One of the mechanism for the generalization is the Weyl squared tensor. Expanding the Weyl squared tensor in the Fefferman-Graham coordinate makes it to be reduced into three terms: divergence of the energy-momentum tensor, shear viscosity and the Weyl anomaly of the boundary theory. Then, the phase transition and the distance between the jumps is controlled by these three terms.
CGR in this way behaves as a function of the generalization parameter and the locations of the jumps. In other words, these two make the CGR to be changed. Considering the behavior of CGR around the phase transition reveals a scaling behavior and critical exponents. It means that the interpretation of the phase transition our of these jumps makes sense. However, CGR satisfies a Callan-Symanzik-like (CSL) equation in field theory in which a beta function is defined in terms of the generalization parameter and an anomalous dimension in terms of the scaling of CGR, which means that the critical behavior of CGR is universal regardless of the generalization parameter.
CSL equation for CGR suggests that in different scale of energy, the minimum required quantum gates is changed. In other words, the change in the scale of energy could change the speed of the processing of information. This recommends that one way to change the speed of the information processing is changing the scale of energy. Suppose that we prepare a quantum computer with a particular speed of processing. One could instead of working on the quantum gates to improve the the speed of processing, work on the scale of energy, put it another way, with a particular array of quantum gates it is enough to change the scale of energy to have fast processing. An interesting application of CSL would be turning the classical computers into a faster one and even faster than our current quantum computer by changing the scale of energy not by the gates or architecture of gates.
Acknowledgment
Author would like to express his gratitude to Monireh Emami for useful discussions.
Appendix A Weyl squared tensor in FG coordinate
The line element in Fefferman-Graham coordinate is given by:
$ds^{2}=G_{\mu\nu}dx^{\mu}dx^{\nu}=\frac{l^{2}}{\rho^{2}}\Big{(}d\rho^{2}+g_{ij}(x,\rho)dx^{i}dx^{j}\Big{)},$ (A-1)
$g_{ij}(x,\rho)=g^{(0)}_{ij}(x)+\rho^{2}g^{(2)}_{ij}(x)+\rho^{4}g^{(4)}_{ij}(x)+\mathcal{O}(\rho^{6}),$ (A-2)
where $g^{(2)}_{ij}=-P_{ij}$ the Schouten tensor of $g^{(0)}_{ij}$ and $g^{(4)}_{ij}\sim<T_{ij}>$ in five dimensions [9]:
$P_{ij}=\frac{1}{2}\Big{(}R_{ij}-\frac{Rg^{(0)}_{ij}}{6}\Big{)},$ (A-3)
where $R_{ij}$ and $R$ are the Ricci tensor and the Ricci scalar of the boundary, respectively. The Weyl tensor, the Riemann tensor and the Ricci tensor by definition could be written, respectively:
$\mathcal{C}_{\mu\nu\rho\sigma}=R_{\mu\nu\rho\sigma}-\Big{(}g_{\mu[\rho}P_{\sigma]\nu}-g_{\nu[\rho}P_{\sigma]\mu}\Big{)},$ (A-4)
$\mathcal{C}_{\rho i\rho j}=R_{\rho i\rho j}-\frac{R_{\rho\rho}}{3}g^{(0)}_{ij}+\mathcal{O}(\rho^{4}),$ (A-5)
$R_{\rho i\rho j}=-\frac{\rho^{2}}{l^{2}}\Big{(}<T_{ij}>-\frac{g^{(0)}_{ij}}{3}<T_{k}^{k}>\Big{)}+\mathcal{O}(\rho^{4}),$ (A-6)
$R_{\rho\rho}=-\frac{4}{l^{2}}+\mathcal{O}(\rho^{2}).$ (A-7)
Then, the Weyl squared tensor in terms of the boundary metric would be:
$\mathcal{C}_{\rho i\rho j}\mathcal{C}^{\rho i\rho j}\sim\Big{(}<T_{ij}>-\frac{g^{(0)}_{ij}}{4}<T_{k}^{k}>\Big{)}^{2}+\mathcal{O}(\rho^{6}).$ (A-8)
In addition, for the other component [9]:
$R_{\rho ijk}\sim\nabla_{i}P_{jk}-\nabla_{j}P_{ik},$ (A-9)
$C_{\rho ijk}C^{\rho ijk}\sim\Big{(}\nabla_{i}P_{jk}-\nabla_{j}P_{ik}\Big{)}^{2}+\mathcal{O}(\rho^{4}),$ (A-10)
$\nabla^{i}g^{(4)}_{ij}=\nabla^{i}P_{ik}P_{j}^{k}-P_{ik}\nabla_{j}P^{ik}.$ (A-11)
The leading term would be:
$\nabla^{i}g^{(4)}_{ij}\sim\nabla^{i}<T_{ij}>\sim\nabla^{i}P_{ij},$ (A-12)
$\mathcal{C}_{\rho ijk}\mathcal{C}^{\rho ijk}\sim\Big{(}\nabla_{i}<T_{jk}>\Big{)}^{2}+\mathcal{O}(\rho^{4})$ (A-13)

### v2

@@
 Abstract
-In this paper, we consider the jumps in the complexity growth rate (CGR) through the complexity=anything proposal.
-It is revealed that bulk fields are in charge of the height and the location of these jumps.
-To put it accurately, their counterparts are divergence of energy momentum tensor, shear viscosity and the Weyl anomaly of the boundary theory.
-The behavior of CGR around these jumps manifests a critical behavior and critical exponents.
-Moreover, it is figured out that CGR satisfies a Callan-Symanzik-like equation around these jumps.
-This equation suggests that the speed of information processing could be changed by the scale of energy.
-Put it another way, by change of the scale of energy one could turn the classical computers into faster processor computers even faster than our current quantum computers.
+Within the ”complexity=anything” proposal of holography, the complexity growth rate (CGR) can exhibit jumps, interpreted as phase transitions.
+We demonstrate that the location and amplitude of these jumps are governed by the dynamics of bulk fields, which, via the fluid-gravity correspondence, map to the boundary energy-momentum tensor.
+The behavior of the CGR near these critical points exhibits scaling and universality.
+We show that the CGR satisfies a Callan-Symanzik-like equation near the transitions.
+We suggest that this enables enhanced computational efficiency by tuning the energy scale.
 1 Introduction
-The AdS/CFT duality establishes a profound correspondence between a gravitational theory in the bulk of anti-de Sitter (AdS) space and a conformal field theory (CFT) on its boundary [1].
-This holographic principle provides a powerful dictionary, allowing quantities that are computationally intractable in one theory—particularly in strongly coupled regimes—to be studied through their dual counterparts in the other.
-A striking example arises in the context of black hole physics, where the growth of Einstein-Rosen bridge volume in the bulk poses a conceptual puzzle: if a black hole is in thermodynamic equilibrium, and its dual boundary theory is likewise equilibrated, what physical meaning should be ascribed to this increasing volume?
-Recent insights suggest that this geometric evolution corresponds to the computational complexity of the boundary theory, specifically the minimal number of fundamental operations required to prepare a target quantum state from a given reference state [2].
-In holography, several conjectures relate bulk geometric quantities to boundary complexity.
-Among these proposals three of them have been widely explored, the complexity=volume (CV) in which the volume of Einstein-Rosen bridges are corresponded to the complexity of the boundary theory, [3], complexity=action (CA), relates the action of the bulk theory restricted to the Wheeler-DeWitt patch to the complexity [4] and complexity=volume (CV2.0) corresponds the spacetime volume of the bulk restricted to the Wheeler-DeWitt patch to the complexity [5].
-More recently, a generalized framework—dubbed complexity=anything—has emerged, positing that arbitrary bulk volumes can encode boundary complexity, offering a broader and more flexible perspective [6].
-Intriguingly, the complexity growth rate (CGR) under this proposal exhibits distinct branches, which some interpretations identify them as signatures of phase transitions [7, 8].
-This work investigates the physical origins of these jumps, addressing two key questions: (1) What determines their heights?
-(2) What controls their temporal locations?
-We demonstrate that the jumps positions are dynamically influenced by bulk fields, which, via the fluid-gravity duality, map to specific boundary quantities: the energy-momentum tensor, Weyl anomaly [9] and shear viscosity [10].
-Consequently, the location of these transitions depends on the divergence of the boundary fluid’s energy-momentum tensor (indicating energy injection), its shear viscosity, and its anomalous conformal behavior.
-Meanwhile, the jump amplitudes are linked to the renormalization group (RG) flow of the boundary theory.
-Remarkably, the CGR transitions exhibit critical phenomena, characterized by scaling behavior and universality.
-Near these critical points, the system obeys a Callan-Symanzik-like equation, where the beta function is expressed in terms of the generalization parameter from the complexity=anything proposal, alongside an anomalous dimension.
-This criticality further displays universal features, independent of the specific choice of the generalized parameter.
-The paper is structured as follows: the section 2 introduces the complexity=anything framework and defines CGR for a generic bulk metric.
-the section 3 elucidates the connection between jump locations and bulk field dynamics.
-The section 4 analyzes the critical behavior of CGR, derives the associated critical exponents, and presents the Callan-Symanzik-like equation in the context of information theory.
-We conclude with a discussion of implications in quantum computers and future directions.
-2 Different branches in $C_{Any}$
+The AdS/CFT correspondence establishes a profound duality between a gravitational theory in anti-de Sitter (AdS) space and a conformal field theory (CFT) on its boundary [1].
+This holographic framework provides a powerful dictionary for studying strongly-coupled quantum systems via their dual classical gravitational descriptions.
+A prominent application lies in black hole physics, where the eternal growth of the Einstein-Rosen bridge volume in the bulk, for a system in thermal equilibrium, posed a conceptual puzzle.
+Recent insights suggest this geometric evolution corresponds to the computational complexity of the boundary quantum state, the minimum number of fundamental operations required to prepare the target state from a reference state [2].
+Several holographic conjectures formalize this connection.
+The most studied are ”complexity=volume” (CV), which identifies complexity with the volume of maximal spatial slices [3], ”complexity=action” (CA), which equates it with the gravitational action on the Wheeler-DeWitt patch [4], and ”complexity=spacetime volume” (CV2.0) [5].
+More recently, a generalized framework, ”complexity=anything”, $\mathcal{C}_{Any}$ has been proposed [6].
+This conjecture posits that complexity can be dual to a vast class of diffeomorphism-invariant observables constructed from arbitrary scalar functionals of the bulk metric.
+A fascinating feature of the $\mathcal{C}_{Any}$ proposal is that the complexity growth rate (CGR) can develop multiple branches and exhibit sudden jumps as a function of time, which have been interpreted as signatures of computational phase transitions in the boundary theory [7, 8].
+This work investigates the physical origin of these transitions.
+We address two central questions:
+  • What boundary physics determines the amplitude of the jumps in the complexity growth rate?
+  • What physics controls their temporal location?
+We find that the jump locations are dynamically set by bulk fields.
+Through the fluid-gravity duality, these map to the boundary energy-momentum tensor [9, 10].
+The jump amplitudes, conversely, are linked to the renormalization group (RG) flow of the boundary theory.
+Remarkably, the CGR near these critical points exhibits scaling behavior and universality.
+We show that it satisfies a Callan-Symanzik-like equation, where the beta function is expressed in terms of the generalization parameter from the $\mathcal{C}_{Any}$, proposal.
+This criticality displays universal features, independent of the specific choice of the generalized functional.
+The paper is organized as follows.
+In the section 2, we review the $\mathcal{C}_{Any}$ framework and derive the CGR for a generic bulk metric.
+The section 3 establishes the connection between the jump locations and the dynamics of boundary field theory quantities.
+In the section 4, we analyze the critical scaling of the CGR, derive the associated critical exponents, and present the central result: a Callan-Symanzik-like equation in the context of information theory.
+We conclude in the section 5 with a discussion of the implications for quantum information processing and future directions.
+2 Complexity growth rate in the $C_{Any}$ proposal
 Recently, $\mathcal{C}_{Any}$ (complexity equals to anything) conjecture has emerged as a generalized version of complexity in the bulk.
@@
 $\mathcal{O}_{F_{1},\Sigma_{F_{2}}}(\Sigma_{CFT})=\frac{1}{G_{N}L}\int_{\Sigma_{F_{2}}}\mathrm{d}^{d}\sigma\sqrt{h}F_{1}(g_{\mu\nu};X^{\mu})$ (1)
-where, $F_{1}$ can be a general scalar function of metric $g_{\mu\nu}$ and an embedding $X^{\mu}(\sigma^{a})$ of the hypersurfaces.
+where $F_{1}$ can be a general scalar function of metric $g_{\mu\nu}$ and an embedding $X^{\mu}(\sigma^{a})$ of the hypersurfaces.
 Furthermore, $\Sigma_{F_{2}}$ is codimension-one hypersurface in the bulk spacetime with boundary time slice $\partial\Sigma_{F_{2}}=\Sigma_{CFT}$.
@@
 $F_{1}=F_{2}=1+\gamma L^{4}\mathcal{C}^{2},$ (4)
-where $\mathcal{C}^{2}$ is the Weyl tensor squared and for $\gamma=0$ the CV is recovered. In the following, we are going to study the $\mathcal{C}_{Any}$ of a generic metric:
+where $\mathcal{C}^{2}$ is the Weyl tensor squared and for $\gamma=0$ the standard CV proposal is recovered. In the following, we are going to study the $\mathcal{C}_{Any}$ of a generic metric:
 $\mathrm{d}s^{2}=g_{tt}\mathrm{d}t^{2}+g_{rr}\mathrm{~d}r^{2}+\frac{r^{2}}{L^{2}}d\Omega^{2}.$ (5)
-To done this, we need to transform the metric to Eddington-Finkelstein coordinates:
+To do this, we need to transform the metric to Eddington-Finkelstein coordinates:
 $\mathrm{d}s^{2}=g_{tt}\mathrm{d}v^{2}-2\sqrt{-g_{tt}g_{rr}}\mathrm{~d}v\mathrm{~d}r+\frac{r^{2}}{L^{2}}d\Omega^{2},$ (6)
@@
 $\mathcal{C}_{Any}(\tau)=\frac{V_{0}}{G_{N}L}\int_{\Sigma}\mathrm{d}\sigma r^{d-1}\sqrt{g_{tt}\dot{v}^{2}-2\sqrt{-g_{tt}g_{rr}}\dot{v}\dot{r}}a(r),$ (8)
-where the dots indicate derivatives with respect to $\sigma$, and $V_{0}$ represents the volume of spatial directions.
+where $a=F_{1}$, the dots indicate derivatives with respect to $\sigma$, and $V_{0}$ represents the volume of spatial directions.
 Supposing $\mathcal{C}_{Any}$ as an action leads to a conserved momentum conjugate to coordinate $v$, because of spacetime is stationary:
@@
 Due to the fact that (8) is diffeomorphism invariant and doesn’t change under reparametrization, the parameter $\sigma$ could be fixed by choosing:
-$\sqrt{g_{tt}\dot{v}^{2}-2\sqrt{-g_{tt}g_{rr}}\dot{v}\dot{r}}=a(r)\sqrt{-g_{tt}g_{rr}}H(r).$ (10)
-Then it is straightforward to derive extremality conditions:
+$\sqrt{g_{tt}\dot{v}^{2}-2\sqrt{-g_{tt}g_{rr}}\dot{v}\dot{r}}=a(r)\sqrt{-g_{tt}g_{rr}}H(r),$ (10)
+where $H$ is an arbitrary function. Then it is straightforward to derive extremality conditions:
 $\dot{r}=\pm\sqrt{\frac{H(r)^{2}P^{2}_{v}}{r^{2(d-1)}}-g_{tt}a(r)^{2}H(r)^{2}},$ (11)
@@
 $\dot{r}^{2}+U(r)=\frac{H(r)^{2}P^{2}_{v}}{r^{2(d-1)}}.$ (14)
-In the symmetric trajectory, the conserved momentum is a function of turning point $r_{min}$, we could redefine the potential as follows:
+In the symmetric trajectory, the conserved momentum is a function of turning point $r_{min}$, we could redefine the potential as follows [^1] At $r_{min}$ the function $H(r)$ is redundant because it is omitted from both sides of (14), so we could redefine new potential which is independent of $H(r)$.:
 $\tilde{U}(r)=g_{tt}a(r)^{2}r^{2(d-1)}.$ (15)
@@
 $=\frac{V_{0}}{G_{N}L}\sqrt{g_{tt}(r_{min})}a(r_{min})r_{min}^{(d-1)}.$ (17)
+A schematic behavior of the complexity growth rate is depicted in Fig.(1).
+Figure 1: a schematic behavior of CGR in terms of the boundary time. The black line shows the phase transition where the complexity has an intersection in terms of $r_{min}$ [7].
 Then, the growth rate of the generalized complexity can be studied by the behavior of the conserved momentum in full-time. The late-time behavior of the growth rate is determined by:
@@
 where $\tilde{r}_{\min}$ is local maximum of the effective potential, and because at this radius time goes to infinity, it is known sometimes as $r_{f}:=\tilde{r}_{\min}$.
-3 What are in charge of these jumps in CGR?
-It has been seen that the complexity equals anything shows some jumps in the complexity growth rate while the former proposals do not.
+3 What controls these jumps in CGR?
+The ”complexity = anything” proposal exhibits jumps in the complexity growth rate while the former proposals do not.
 This phenomenon raises the question that why there are these jumps in this proposal and what they refer to.
-Are they related to a physical phenomena?
-In this section we answer to the questions affirmatively and seeking the reason on the boundary theory of the bulk.
-The jump occurs around the asymptotic behavior of CGR, the extrimum values of $P_{v}$, when the potential (15) has a local maxima, in other words $\frac{d\tilde{U}}{dr}=0$.
+Are they related to a physical phenomenon?
+In this section we answer to the questions affirmatively and seek the reason in the boundary theory of the bulk.
+The jump occurs around the asymptotic behavior of CGR, the extremum values of $P_{v}$, when the potential (15) has a local maxima, in other words $\frac{d\tilde{U}}{dr}=0$ around the jumps.
+A schematic behavior of the potential $\tilde{U}$ is demonstrated in Fig.(2).
+Figure 2: A schematic behavior of the potential $\tilde{U}$ in terms of radial coordinate $r$ and the horizon $R$.
+Since the jump locations $r_{i}$ are determined by the behavior of $\tilde{U}(r)$, we can express this condition using the Fefferman-Graham expansion at finite $r_{i}$ cut-off, which directly links it to the boundary energy-momentum tensor and its properties.
 For a generic solution we have got:
@@
 $g_{tt}a^{2}r^{2(d-1)}=cte:=\alpha.$ (21)
-In the following we set $\alpha=1$.
-By rescaling $g_{tt}\rightarrow r^{-2(d-1)-2}g_{tt}=g_{t^{\prime}t^{\prime}}$ and $dt^{2}\rightarrow r^{2(d-1)+2}dt^{2}=dt^{\prime 2}$ one could rewrite (21) as following:
-$r^{-2}g_{tt}a^{2}$ $=1$ (22)
-$r^{-2}g_{tt}(1+\gamma C^{2})^{2}$ $=1.$ (23)
-At the leading term in terms of $\gamma$ [^1] We suppose that $\gamma\ll 1$.:
-$r^{-2}g_{tt}+2\gamma r^{-2}g_{tt}C^{2}$ $=1$ (24)
+In the following we set $\alpha=1$. One could rewrite (21) in Fefferman-Graham coordinates as following (at small $\gamma\ll 1$):
+$r^{2(d-1)}\Big(r^{2}g_{tt}^{(0)}+g_{tt}^{(2)}+\frac{g_{tt}^{(4)}}{r^{2}}+...\Big)\Big(1+2\gamma\mathcal{C}^{2}\Big)=1.$ (22)
 If one writes the bulk Weyl squared tensor in terms of the boundary coordinates:
-$\mathcal{C}_{\mu\nu\rho\sigma}\mathcal{C}^{\mu\nu\rho\sigma}=\mathcal{C}_{ijkl}\mathcal{C}^{ijkl}+4\mathcal{C}_{i\rho j\rho}\mathcal{C}^{i\rho j\rho}+4\mathcal{C}_{ijk\rho}\mathcal{C}^{ijk\rho},$ (25)
-where ijkijk...
-are boundary coordinates and $\rho$ the radial.
-Using the Fefferman-Graham coordinate to write the bulk Weyl squared tensor in terms of the boundary values it is given by (see Appendix A)  [^2] It should be noted that we used the FG coordinates at the finite cut-off $r=r_{j}$ where $r_{j}$ is a location of a local maxima of the potential.:
-$C^{2}$ $=\mathscr{C}+\Big{(}<T_{ij}>-\frac{g_{ij}^{(0)}}{4}<T_{k}^{k}>\Big{)}^{2}+\Big{(}\nabla_{i}<T_{jk}>\Big{)}^{2}+\mathcal{O}(\rho^{2}),$ (26)
-$=C_{(0)}^{2}+\mathcal{O}(\rho^{2}),$ (27)
-where $\mathscr{C}^{2}$ is the boundary Weyl tensor squared, $T_{ij}>$ the expectation value of the boundary energy-momentum tensor and $g_{ij}^{(0)}$ the induced metric on the boundary.
-The first term in the right-hand-side of (27) is the Weyl anomaly of the boundary theory [9], the second term is related to the shear viscosity of the boundary theory [10] and the third term is not conserved in the presence of matter that shows the flow of the matter through the boundary [11].
-By rewriting (24) in the FG coordinates one gets [^3] The coordinate change in (23) makes the following formula, (28), to be written in terms of the first order of the Weyl squared tensor expansion.
-If one does not use the coordinate change, (28) is written in terms of the other orders of the Weyl squared tensor expansion.
-However, we know that the other terms could be rewritten in terms of the first order.
-Then, the coordinate change makes the result more clear.
-In addition, we should note that $r_{i}$ appears due to the FG coordinates at $r_{i}$ cut-off:
-$g^{(0)}_{tt}+2\gamma g^{(0)}_{tt}\Bigg{(}\mathscr{C}+\Big{(}<T_{ij}>-\frac{g_{ij}^{(0)}}{4}<T_{k}^{k}>\Big{)}^{2}+\Big{(}\nabla_{i}<T_{jk}>\Big{)}^{2}+\mathcal{O}(\rho^{2})\Bigg{)}=1.$ (28)
-(28) shows that the jumps in the complexity are refereed to the following quantities in the boundary theory:
-  • The Weyl anomaly, $\mathscr{C}$.
-  • The shear viscosity, $\Big{(}<T_{ij}>-\frac{g_{ij}^{(0)}}{4}<T_{k}^{k}>\Big{)}^{2}$.
-  • Non-conservation of the energy-momentum on the boundary, $\Big{(}\nabla_{i}<T_{jk}>\Big{)}^{2}$.
-  • The boundary metric, $g^{(0)}_{tt}$.
-  • The parameter of the generalization, $\gamma$.
-If there is no Weyl anomaly in the boundary theory and no shear viscosity, then it is expected that the jumps in the complexity are related to:
-$C_{(0)}^{2}=\Big{(}\nabla_{i}<T_{jk}>\Big{)}^{2}=\frac{1-g^{(0)}_{tt}}{2\gamma g^{(0)}_{tt}}$. (29)
-In [11] it has been shown that the non-conservation of the energy-momentum tensor on the boundary is related to the presence of the derivative of a bulk field, say a scalar field:
-$\nabla^{i}<T_{ij}>=<\hat{O}>\partial_{j}\phi^{(0)},$ (30)
-where $\mathcal{O}$ is the operator sourced by the bulk field $\phi$.
-It has been revealed that the value of the bulk fields on the boundary could determine the location of the jumps in CGR.
-Having done that, there would be a question that if the distance of the jumps is related to a physical quantity or not?
+$\mathcal{C}^{2}=\mathcal{C}_{\mu\nu\rho\sigma}\mathcal{C}^{\mu\nu\rho\sigma}=\mathcal{C}_{ijkl}\mathcal{C}^{ijkl}+4\mathcal{C}_{i\rho j\rho}\mathcal{C}^{i\rho j\rho}+4\mathcal{C}_{ijk\rho}\mathcal{C}^{ijk\rho}+$..., (23)
+where $i,j,k,l$ are boundary coordinates and $\rho=\frac{1}{r}$ the radial.
+Using the Fefferman-Graham coordinate to write the bulk Weyl squared tensor in terms of the boundary values it is given by (see Appendix A)  [^2] It should be noted that we used the FG coordinates at the finite cut-off $r=r_{j}$ where $r_{j}$ is a location of a local maxima of the potential.
+In addition, (24) is the Wey squared tensor in $d+1=4+1$ dimensions.:
+$\mathcal{C}^{2}=\rho^{4}\mathscr{C}^{2}+\rho^{4}Log\rho^{2}\mathcal{A}+\rho^{6}\mathcal{B}+\rho^{8}\Big((\nabla_{i}g^{(4)}_{jk})^{2}+g^{(4)}_{ij}g^{(4)ij}\Big)+$..., (24)
+$=\rho^{4}\mathcal{C}_{(4)}^{2}+\mathcal{O}(\rho^{6}),$ (25)
+where $\mathscr{C}^{2}$ is the boundary Weyl tensor squared, $\mathcal{A}$ and $\mathcal{B}$ are constant [11].
+(24) is a pure geometrical expansion of the bulk Weyl squared tensor.
+Expressing (22) in the FG coordinates at order $\rho^{2}$ yields:
+$\rho^{2}g^{(4)}_{tt}+2\gamma\frac{g^{(0)}_{tt}}{\rho^{2}}\Bigg(\rho^{4}\mathscr{C}^{2}+\rho^{4}Log\rho^{2}\mathcal{A}+\mathcal{O}(\rho^{4})\Bigg)=1.$ (26)
+(26) shows that the jumps in the complexity are governed by dynamics of the energy-momentum tensor of the boundary  [^3] The expectation value of the boundary energy-momentum tensor is computed by the Brown-York Lagrangian (BY) in the bulk $T_{ij}>=\frac{\partial S_{ren.}}{\partial g^{(0)}_{ij}}$.
+BY term expansion contains $g^{(4)}_{ij}$ [9].
+In addition, the Weyl anomaly of the boundary theory is computed by the variation of the bulk Lagrangian under the conformal transformation.
+The density of the holographic Weyl anomaly comes from the coefficient of logarithm divergence [9].
+Moreover, the Weyl anomaly in the boundary theory comes from the correlation of the boundary energy-momentum tensor.
+All in all, the bulk Weyl squared tensor is governed by the boundary energy-momentum tensor..
+If there is non-vanishing Weyl squared tensor at the boundary, then it is expected that the jumps in the complexity at the finite cut-off $\rho=\rho_{i}$ are related to 444(27) is written in general dimension, where the logarithmic term is of order $d$.:
+$\mathscr{C}^{2}=\frac{\rho_{i}^{-2}-g^{(4)}_{tt}}{2\gamma g^{(0)}_{tt}}$. (27)
+The above relation raises the question of whether the distance between the jumps corresponds to a physical quantity.
 The distance between the jumps could be realized by $d_{jumps}=|r_{i}-r_{j}|$ where $r_{i}$ means the location where the ith"i"th jump occurs at.
@@
 In other words, roots of (21).
-The distance between roots depends on the quantities in (21), quantities in $g_{tt}$ and $\gamma$ the coefficient of the generalization:
-$g_{tt}a^{2}r^{-2}=1$ (31)
-$Q(r_{i})=g_{tt}a^{2}r^{-2}-1$ (32)
-where $r_{i}$ is the ith"i"th root of (21). The dependence could be regarded as:
-$\frac{\partial r_{i}}{\partial a}=\frac{\frac{\partial Q}{\partial a}}{\frac{\partial Q}{\partial r}}\Big{|}_{r=r_{i}}=\frac{\partial Q}{\partial a}\Big{|}_{r=r_{i}}=\frac{\partial\tilde{U}}{\partial a}\Big{|}_{r=r_{i}}$ (33)
-where $\frac{\partial r_{i}}{\partial a}$ determines the dependency of the distance to $a$, one the quantities in $g_{tt}$ and $\gamma$.
-As an example we show that how jumps are controlled by the parameter of the generalization:
-$\frac{\partial r_{i}}{\partial\gamma}=2\Big{(}g_{tt}^{(0)}C_{(0)}^{2}\Big{)}\Big{|}_{r=r_{i}}.$ (34)
+The distance between roots depends on the quantities in (21), quantities in $g_{tt}$ and $\gamma$ the coefficient of the generalization.
+As an example we show that how jumps are controlled by the parameter of the generalization by (21):
+$\frac{\partial r_{i}}{\partial\gamma}=-\frac{\mathcal{C}^{2}}{(d-1)r}\Big|_{r=r_{i}}.$ (28)
 As a consequence the distance between jumps could be written:
-$\frac{\partial\big{(}r_{i}-r_{j}\big{)}}{\partial\gamma}=2\Bigg{(}g_{tt}^{(0)}C_{(0)}^{2}\Big{|}_{r=r_{i}}-g_{tt}^{(0)}C_{(0)}^{2}\Big{|}_{r=r_{j}}\Bigg{)},$ (35)
+$\frac{\partial\big(r_{i}-r_{j}\big)}{\partial\gamma}=\frac{1}{1-d}\Bigg(\frac{\mathcal{C}^{2}}{r}\Big|_{r=r_{i}}-\frac{\mathcal{C}^{2}}{r}\Big|_{r=r_{j}}\Bigg),$ (29)
 which means that the distance between jumps is encoded in the RG flow of the boundary theory i.e.
-the divergence of the energy-momentum tensor, shear viscosity and the Weyl anomaly at different energy scales.
+the dynamics of the energy-momentum tensor of the boundary at different energy scales.
 Moreover, (29) tells about the scale-dependent origin of the distance of the jumps in CGR.
-4 Critical exponents
-As it has been revealed that the location of these jumps are scale dependent.
+4 Critical behavior and a Callan-Symanzik-like equation
+As shown above, the locations of these jumps are scale dependent.
 This result suggests that the CGR itself could behave differently in various scales.
 In this section we are going to study the behavior of CGR in different scales.
-In this manner CGR is a function of the generalization parameter; however, the location of these jumps $r_{i}$ is a function of the parameter, so we could use a chain rule:
+In this manner CGR is a function of the generalization parameter; however, the location of these jumps $r_{i}$ is a function of the parameter, so we could use a chain rule and (28):
 $\frac{d\dot{\mathcal{C}}}{d\gamma}\Big|_{r_{i}}=$ $\frac{\partial\dot{\mathcal{C}}}{\partial\gamma}\Big|_{r_{i}}+\frac{\partial\dot{\mathcal{C}}}{\partial r}\frac{\partial r}{\partial\gamma}\Big|_{r_{i}}$ (30)
-$=$ $\frac{\partial\dot{\mathcal{C}}}{\partial\gamma}\Big{|}_{r_{i}}+2g_{tt}^{(0)}C_{(0)}^{2}\frac{\partial\dot{\mathcal{C}}}{\partial r}\Big{|}_{r_{i}}.$ (37)
-The above result appears when we are working in $t^{\prime}$ coordinate and Fefferman-Graham coordinate at $r_{i}$ cut-off. One could proceed with $t$ coordinate and $r_{i}$ cut-off, then:
-$\frac{d\dot{\mathcal{C}}}{d\gamma}\Big{|}_{r_{i}}=\frac{\partial\dot{\mathcal{C}}}{\partial\gamma}\Big{|}_{r_{i}}+2r_{i}^{2(d)}g_{tt}^{(0)}C_{(0)}^{2}\frac{\partial\dot{\mathcal{C}}}{\partial r}\Big{|}_{r_{i}}.$ (38)
+$=$ $\frac{\partial\dot{\mathcal{C}}}{\partial\gamma}\Big|_{r_{i}}+\frac{\mathcal{C}^{2}}{(1-d)r}\frac{\partial\dot{\mathcal{C}}}{\partial r}\Big|_{r_{i}}.$ (31)
+One could proceed with the leading order:
+$\frac{d\dot{\mathcal{C}}}{d\gamma}\Big|_{r_{i}}=\frac{\partial\dot{\mathcal{C}}}{\partial\gamma}\Big|_{r_{i}}+\frac{\mathscr{C}^{2}}{(1-d)r^{5}_{i}}\frac{\partial\dot{\mathcal{C}}}{\partial r}\Big|_{r_{i}}.$ (32)
 Around $r_{i}$, where it is a late time and asymptotic behavior of $\dot{\mathcal{C}}$ in Fefferman-Graham coordinates at $r_{i}$ cut-off, CGR at the leading order could be written as:
 $\dot{\mathcal{C}}\sim P_{v}=\sqrt{\tilde{U}}=\sqrt{g_{tt}a^{2}r^{2(d-1)}}$ (33)
-$=\Bigg{(}r_{i}^{2(d-1)}\Big{(}r_{i}^{2}g_{tt}^{(0)}+g_{tt}^{(1)}+\frac{g_{tt}^{(2)}}{r_{i}^{2}}+...\Big{)}\Big{(}C_{(0)}^{2}+\frac{C_{(1)}^{2}}{r_{i}^{2}}+...\Big{)}\Bigg{)}^{\frac{1}{2}}$ (40)
-$\approx\Big{(}r_{i}^{2d}g_{tt}^{(0)}+r_{i}^{2d}(2\gamma g_{tt}^{(0)}C_{(0)}^{2})\Big{)}^{\frac{1}{2}}.$ (41)
+$=\Bigg(r_{i}^{2(d-1)}\Big(r_{i}^{2}g_{tt}^{(0)}+g_{tt}^{(2)}+\frac{g_{tt}^{(4)}}{r_{i}^{2}}+...\Big)\Big(1+\gamma(\frac{\mathcal{C}_{(4)}^{2}}{r_{i}^{4}}+\frac{\mathcal{C}_{(6)}^{2}}{r_{i}^{6}}+...)\Big)^{2}\Bigg)^{\frac{1}{2}}$ (34)
+$=\Big(r_{i}^{2d}g_{tt}^{(0)}+r_{i}^{2(d-1)}g_{tt}^{(2)}+r_{i}^{2(d-2)}(g_{tt}^{(4)}+2\gamma g_{tt}^{(0)}\mathcal{C}_{(4)}^{2})+r_{i}^{2(d-3)}(g_{tt}^{(5)}+2\gamma g_{tt}^{(4)}\mathcal{C}_{(4)}^{2})$
+$+r_{i}^{2(d-4)}(\gamma^{2}g_{tt}^{(0)}\mathcal{C}_{(4)}^{2}\mathcal{C}_{(4)}^{2})+...\Big)^{\frac{1}{2}}.$ (35)
 As a consequence, if $\gamma\rightarrow\lambda^{\nu}\gamma$ and $r_{i}\rightarrow\lambda^{\mu}r_{i}$, then $\dot{\mathcal{C}}(\lambda^{\mu}r_{i},\lambda^{\nu}\gamma)=\lambda^{\Delta}\dot{\mathcal{C}}(r_{i},\gamma)$, where $\Delta$, $\mu$ and $\nu$ are critical exponents.
@@
 Combining (37) and (38):
-$\Big{(}(2r_{i}^{2d}\nu\gamma g_{tt}^{(0)}C_{(0)}^{2}-\mu r_{i})\frac{\partial}{\partial r_{i}}-\nu\gamma\frac{d}{d\gamma}+\Delta\Big{)}\dot{\mathcal{C}}=0,$ (45)
-where $(2r^{2d}\nu\gamma g_{tt}^{(0)}C_{(0)}^{2}-\mu r_{i})\frac{\partial\dot{\mathcal{C}}}{\partial r}$ shows how CGR is changed by the scale of the energy $r_{i}$, $\nu\gamma\frac{d\dot{\mathcal{C}}}{d\gamma}$ indicating how CGR is changed by the generalization parameter, in other words, if the CGR is increasing by increase in $\gamma$ or decreasing.
+$\Big((\frac{\nu\gamma\mathscr{C}^{2}}{(1-d)r_{i}^{5}}-\mu r_{i})\frac{\partial}{\partial r_{i}}-\nu\gamma\frac{d}{d\gamma}+\Delta\Big)\dot{\mathcal{C}}=0,$ (39)
+where $(\frac{\nu\gamma\mathscr{C}^{2}}{(1-d)r_{i}^{5}}-\mu r_{i})\frac{\partial}{\partial r_{i}}$ shows how CGR is changed by the scale of the energy $r_{i}$, $\nu\gamma\frac{d\dot{\mathcal{C}}}{d\gamma}$ indicating how CGR is changed by the generalization parameter, in other words, if the CGR is increasing by increase in $\gamma$ or decreasing.
 We call (47) Callan-Symanzik-like equation in information theory, where $r_{i}$ is the information renormalization scale, $\beta=-\nu\gamma$ beta function in information theory and $\Delta$ information anomalous dimension.
-(41) suggests that CGR near $r_{i}$ is scaled:
-$\dot{\mathcal{C}}(\lambda^{\mu}r_{i},\lambda^{\nu}\gamma)$ $\sim\Big{(}\lambda^{2d\mu}r_{i}^{2d}g_{tt}^{(0)}+\lambda^{2d\mu+\nu}r_{i}^{2d}(2\gamma g_{tt}^{(0)}C_{(0)}^{2})\Big{)}^{\frac{1}{2}}$ (46)
+(35) suggests that CGR at $\gamma\ll 1$, near $r_{i}$ is scaled:
+$\dot{\mathcal{C}}(\lambda^{\mu}r_{i},\lambda^{\nu}\gamma)\sim\Big(\lambda^{2d\mu}r_{i}^{2d}g_{tt}^{(0)}+\lambda^{2(d-1)\mu}r_{i}^{2(d-1)}g_{tt}^{(2)}+\lambda^{2(d-2)\mu}r_{i}^{2(d-2)}g_{tt}^{(4)}$
+$+\lambda^{2(d-2)\mu+\nu}r_{i}^{2(d-2)}2\gamma g_{tt}^{(0)}\mathcal{C}_{(4)}^{2}+\lambda^{2(d-3)\mu}r_{i}^{2(d-3)}g_{tt}^{(5)}+\lambda^{2(d-3)\mu+\nu}r_{i}^{2(d-3)}2\gamma g_{tt}^{(4)}\mathcal{C}_{(4)}^{2}\Big)^{\frac{1}{2}}$ (40)
 $=\lambda^{\Delta}\dot{\mathcal{C}}(r_{i},\gamma)$ (41)
-Now we could consider the validity of critical exponents:
-$\Delta=d\mu$ (48)
-$\Delta=d\mu+\frac{\nu}{2}$ (49)
-The solution for the critical exponents would be $\nu=0$ and $\Delta=d\mu$. This solution demonstrates a universal critical behavior around the $r_{i}$’s or jumps of CGR.
-However, for large $\gamma$ we could write $a^{2}\approx 1+\gamma^{2}C^{4}$. Then (41) is rewritten as:
-$\dot{\mathcal{C}}\sim P_{v}=\sqrt{g_{tt}a^{2}r^{2(d-1)}}=\Big{(}r_{i}^{2d}g_{tt}^{(0)}+r_{i}^{2d}(\gamma^{2}g_{tt}^{(0)}(C_{(0)}^{2})^{2})\Big{)}^{\frac{1}{2}}.$ (51)
-The critical exponents would be:
-$\Delta=d\mu$ (52)
-$\Delta=d\mu+\nu$ (53)
-For the large $\gamma$ the same solution appears.
-This result illustrates a universal critical behavior around the phase transition of CGR regardless of the generalization parameter $\gamma$.
-Looking again at (55) by the solution reveals:
-$\Big{(}-\mu r_{i}\frac{\partial}{\partial r_{i}}+d\mu\Big{)}\dot{\mathcal{C}}=0,$ (55)
+Now we could consider the validity of critical exponents for arbitrary dimensions:
+$\Delta=\frac{\nu}{2},$ (42)
+$\mu=0.$ (43)
+However, for large $\gamma\gg 1$ we could write $a^{2}\approx 1+\gamma^{2}C^{4}$. Then (35) is scaled as:
+$\dot{\mathcal{C}}\sim P_{v}=\sqrt{g_{tt}a^{2}r^{2(d-1)}}=\Big(\lambda^{2d\mu}r_{i}^{2d}g_{tt}^{(0)}+\lambda^{2(d-1)\mu}r_{i}^{2(d-1)}g_{tt}^{(2)}+\lambda^{2(d-2)\mu}r_{i}^{2(d-2)}g_{tt}^{(4)}$
+$+\lambda^{2(d-3)\mu}r_{i}^{2(d-3)}g_{tt}^{(5)}+\lambda^{2(d-4)\mu+2\nu}r_{i}^{2(d-4)}\gamma^{2}g_{tt}^{(0)}\mathcal{C}_{(4)}^{2}\mathcal{C}_{(4)}^{2}\Big)^{\frac{1}{2}}.$ (44)
+The critical exponents for arbitrary dimensions would be:
+$\Delta=\nu,$ (45)
+$\mu=0.$ (46)
+As (42) and (45) show, the scaling of CGR is different at different regimes of the generalization parameter. Looking again at (47) by the solution reveals:
+$\Big((\frac{\nu\gamma\mathscr{C}^{2}}{(1-d)r_{i}^{5}})\frac{\partial}{\partial r_{i}}-\nu\gamma\frac{d}{d\gamma}+\Delta\Big)\dot{\mathcal{C}}=0,$ (47)
 The solution of the above equation would be:
-$\dot{\mathcal{C}}\sim r_{i}^{d}(\gamma)+\alpha,$ (56)
-where $\alpha$ is a constant in (21). By the fact that $\dot{\mathcal{C}}$ is constant at $r_{i}$, then it could be written as:
-$\dot{\mathcal{C}}\sim\Big{(}r-r_{i}(\gamma)\Big{)}^{d}+\alpha.$ (57)
-This solution shows that the critical behavior of CGR around the critical point $r_{i}$. Then, the critical exponents $\Delta=d\mu=d$.
+$\dot{\mathcal{C}}=\gamma^{n}\mathcal{F}(r^{6})+\alpha,$ (48)
+where $\mathcal{F}$ is a function of $r^{6}$, $n=1$ at $\gamma\gg 1$, $n=\frac{1}{2}$ at $\gamma\ll 1$ and $\alpha$ is the constant in (21).
+By the fact that $\dot{\mathcal{C}}$ is constant at $r_{i}$, then it could be written as:
+$\dot{\mathcal{C}}=\gamma^{n}\mathcal{F}(r^{6}-r_{i}^{6}(\gamma)+\alpha.$ (49)
+This indicates how the CGR scales with the correlation length $|r-r_{i}|$ near the critical point for any arbitrary dimension.
 5 Conclusion
-The AdS/CFT by corresponding a gravitational theory to a quantum theory could alleviate the perturbation calculation in strongly-coupled systems.
+The AdS/CFT correspondence, by relating a gravitational theory to a quantum field theory, can alleviate the difficulties of perturbation theory in strongly-coupled systems.
 Exploring the counterpart quantities in the AdS/CFT make it a strong technique to study the physical theories.
-Computational complexity as one of these quantities has been considered in holography and provided different proposal.
-Recently, a proposal known as complexity equals anything has been introduced in which the volume of a generalized surface is corresponded to the complexity of the boundary theory.
+Computational complexity, as one of these quantities, has been extensively studied in holography through several proposals.
+Recently, a proposal known as complexity equals anything has been introduced where the complexity is identified with the volume of a generalized hypersurface.
 Studying the complexity of various models through this proposal shows some branches in CGR.
 Moreover, these branches has been considered as a phase transition in CGR.
-Because of the fact that the phase transition appeared in complexity equals anything, this suggests that the generalized surfaces could make the transition.
-One of the mechanism for the generalization is the Weyl squared tensor.
-Expanding the Weyl squared tensor in the Fefferman-Graham coordinate makes it to be reduced into three terms: divergence of the energy-momentum tensor, shear viscosity and the Weyl anomaly of the boundary theory.
-Then, the phase transition and the distance between the jumps is controlled by these three terms.
+Since phase transitions appear in the ”complexity = anything” framework, this suggests that the generalized surfaces could make the transition.
+One mechanism for the generalization is the Weyl squared tensor.
+Expanding the Weyl squared tensor in the Fefferman-Graham coordinate makes it to be reduced into two terms: divergence of the energy-momentum tensor and the Weyl squared tensor of the boundary which is only a geometrical expansion.
+However, the phase transitions and jump separations are governed by the boundary energy-momentum tensor.
 CGR in this way behaves as a function of the generalization parameter and the locations of the jumps.
-In other words, these two make the CGR to be changed.
+In other words, these two parameters control CGR.
 Considering the behavior of CGR around the phase transition reveals a scaling behavior and critical exponents.
-It means that the interpretation of the phase transition our of these jumps makes sense.
-However, CGR satisfies a Callan-Symanzik-like (CSL) equation in field theory in which a beta function is defined in terms of the generalization parameter and an anomalous dimension in terms of the scaling of CGR, which means that the critical behavior of CGR is universal regardless of the generalization parameter.
+This shows that interpreting these jumps as phase transitions is meaningful.
+However, CGR satisfies a Callan-Symanzik-like (CSL) equation in field theory in which a beta function is defined in terms of the generalization parameter and an anomalous dimension in terms of the scaling of CGR, which means that the critical behavior of CGR is universal regardless of the dimension.
 CSL equation for CGR suggests that in different scale of energy, the minimum required quantum gates is changed.
 In other words, the change in the scale of energy could change the speed of the processing of information.
-This recommends that one way to change the speed of the information processing is changing the scale of energy.
+This may suggest that modifying the energy scale could provide a means to control information processing speed.
 Suppose that we prepare a quantum computer with a particular speed of processing.
-One could instead of working on the quantum gates to improve the the speed of processing, work on the scale of energy, put it another way, with a particular array of quantum gates it is enough to change the scale of energy to have fast processing.
-An interesting application of CSL would be turning the classical computers into a faster one and even faster than our current quantum computer by changing the scale of energy not by the gates or architecture of gates.
+One could instead of working on the quantum gates to improve the the speed of processing, modify the energy scale to achieve faster processing.
+As a future work one could use the application of CSL to evaluate how it improves the processing in real quantum computers.
 Acknowledgment
-Author would like to express his gratitude to Monireh Emami for useful discussions.
+The authors would like to express their gratitude to Monireh Emami for useful discussions. In addition, OpenAI’s ChatGPT has been used to enhance the clarity of this manuscript.
 Appendix A Weyl squared tensor in FG coordinate
-The line element in Fefferman-Graham coordinate is given by:
+In this section the derivation of (24) is provided. The line element in Fefferman-Graham coordinate is given by:
 $ds^{2}=G_{\mu\nu}dx^{\mu}dx^{\nu}=\frac{l^{2}}{\rho^{2}}\Big(d\rho^{2}+g_{ij}(x,\rho)dx^{i}dx^{j}\Big),$ (A-1)
-$g_{ij}(x,\rho)=g^{(0)}_{ij}(x)+\rho^{2}g^{(2)}_{ij}(x)+\rho^{4}g^{(4)}_{ij}(x)+\mathcal{O}(\rho^{6}),$ (A-2)
+$g_{ij}(x,\rho)=g^{(0)}_{ij}(x)+\rho^{2}g^{(2)}_{ij}(x)+\rho^{4}g^{(4)}_{ij}(x)+...+\rho^{d}(g^{(d)}_{ij}+h^{(d)}_{ij}Log\rho^{2})+.$... (A-2)
 where $g^{(2)}_{ij}=-P_{ij}$ the Schouten tensor of $g^{(0)}_{ij}$ and $g^{(4)}_{ij}\sim<T_{ij}>$ in five dimensions [9]:
 $P_{ij}=\frac{1}{2}\Big(R_{ij}-\frac{Rg^{(0)}_{ij}}{6}\Big),$ (A-3)
-where $R_{ij}$ and $R$ are the Ricci tensor and the Ricci scalar of the boundary, respectively. The Weyl tensor, the Riemann tensor and the Ricci tensor by definition could be written, respectively:
-$\mathcal{C}_{\mu\nu\rho\sigma}=R_{\mu\nu\rho\sigma}-\Big{(}g_{\mu[\rho}P_{\sigma]\nu}-g_{\nu[\rho}P_{\sigma]\mu}\Big{)},$ (A-4)
-$\mathcal{C}_{\rho i\rho j}=R_{\rho i\rho j}-\frac{R_{\rho\rho}}{3}g^{(0)}_{ij}+\mathcal{O}(\rho^{4}),$ (A-5)
-$R_{\rho i\rho j}=-\frac{\rho^{2}}{l^{2}}\Big{(}<T_{ij}>-\frac{g^{(0)}_{ij}}{3}<T_{k}^{k}>\Big{)}+\mathcal{O}(\rho^{4}),$ (A-6)
-$R_{\rho\rho}=-\frac{4}{l^{2}}+\mathcal{O}(\rho^{2}).$ (A-7)
-Then, the Weyl squared tensor in terms of the boundary metric would be:
-$\mathcal{C}_{\rho i\rho j}\mathcal{C}^{\rho i\rho j}\sim\Big{(}<T_{ij}>-\frac{g^{(0)}_{ij}}{4}<T_{k}^{k}>\Big{)}^{2}+\mathcal{O}(\rho^{6}).$ (A-8)
-In addition, for the other component [9]:
-$R_{\rho ijk}\sim\nabla_{i}P_{jk}-\nabla_{j}P_{ik},$ (A-9)
-$C_{\rho ijk}C^{\rho ijk}\sim\Big{(}\nabla_{i}P_{jk}-\nabla_{j}P_{ik}\Big{)}^{2}+\mathcal{O}(\rho^{4}),$ (A-10)
-$\nabla^{i}g^{(4)}_{ij}=\nabla^{i}P_{ik}P_{j}^{k}-P_{ik}\nabla_{j}P^{ik}.$ (A-11)
-The leading term would be:
-$\nabla^{i}g^{(4)}_{ij}\sim\nabla^{i}<T_{ij}>\sim\nabla^{i}P_{ij},$ (A-12)
-$\mathcal{C}_{\rho ijk}\mathcal{C}^{\rho ijk}\sim\Big{(}\nabla_{i}<T_{jk}>\Big{)}^{2}+\mathcal{O}(\rho^{4})$ (A-13)
+In $4+1$ dimensions (d=4) the expansion would be:
+$g_{ij}(x,\rho)=g^{(0)}_{ij}(x)+\rho^{2}g^{(2)}_{ij}(x)+\rho^{4}\Big(g^{(4)}_{ij}(x)+h^{(4)}_{ij}(x)Log\rho^{2}\Big)+\mathcal{O}(\rho^{6}).$ (A-4)
+Then, the Riemann tensor is given by:
+$\Gamma_{\rho\rho}^{\rho}=-\frac{1}{\rho},$ (A-5)
+$\Gamma_{ij}^{\rho}=-\frac{1}{\rho}g_{ij}+\frac{1}{2}\partial_{\rho}g_{ij}=2\rho g^{(2)}_{ij}+4\rho^{3}(g^{(4)}_{ij}+h^{(4)}_{ij}Log\rho^{2})+$..., (A-6)
+$\Gamma_{\rho j}^{i}=-\frac{1}{\rho}\delta_{j}^{i}+\frac{1}{2}g^{ik}\partial_{\rho}g_{kj}$ (A-7)
+$R_{ijkl}(G)=R_{ijkl}(g|_{\rho})-\frac{1}{\rho^{2}}(g_{ik}g_{jl}-g_{il}g_{jk})+\frac{1}{4}\partial_{\rho}g_{ik}\partial_{\rho}g_{jl}-\frac{1}{4}\partial_{\rho}g_{il}g_{jk}+$... (A-8)
+where $g|_{\rho}=g^{(0)}_{ij}+\rho^{2}g^{(2)}_{ij}$+... . For the other component:
+$R_{\rho i\rho j}(G)=-\frac{1}{2}\partial_{\rho}^{2}g_{ij}+\frac{1}{4}g^{kl}\partial_{\rho}g_{ik}\partial_{\rho}g_{jl}=-\frac{1}{\rho^{2}}g^{(0)}_{ij}$+.... (A-9)
+The other component goes as:
+$R_{\rho ijk}(G)=\frac{1}{2}(\nabla_{j}\partial_{\rho}g_{ik}-\nabla_{k}\partial_{\rho}g_{ij})=\nabla_{j}(\rho g^{(2)}_{ik})-\nabla_{k}(\rho g^{(2)}_{ij})$. (A-10)
+where $\nabla$ is with respect to $g|_{\rho}$. The Ricci tensor would be:
+$R_{\rho\rho}=-\frac{1}{2}g^{ij}\partial_{\rho}^{2}g_{ij}+\frac{1}{4}g^{ik}g^{jl}\partial_{\rho}g_{ij}\partial_{\rho}g_{kl},$ (A-11)
+$R_{\rho i}=\frac{1}{2}g^{jk}(\nabla_{k}\partial_{\rho}g_{ij}-\nabla_{i}\partial_{\rho}g_{jk}),$ (A-12)
+$R_{ij}=R_{ij}(g|_{\rho})-\frac{1}{2}\partial_{\rho}^{2}g_{ij}+\frac{1}{4}g^{kl}\partial_{\rho}g_{ik}\partial_{\rho}g_{jl}+...=R_{ij}(g|_{\rho})-4g^{(2)}_{ij}$+.... (A-13)
+The Ricci scalar would be:
+$R=-\frac{20}{l^{2}}+\rho^{2}R(g^{(0)})+.$... (A-14)
+The Weyl tensor is given by:
+$\mathcal{C}_{ijkl}(G)=\mathcal{C}_{ijkl}(g^{(0)})+\rho^{2}g^{(2)}+\rho^{4}(g^{(4)}+h^{(4)}Log\rho^{2})+.$... (A-15)
+Now, the Weyl squared tenor is computed as:
+$\mathcal{C}^{2}=\rho^{4}\mathscr{C}+\rho^{4}Log\rho^{2}\mathcal{A}+\rho^{6}\mathcal{B}+\rho^{8}\Big((\nabla_{i}g^{(4)}_{jk})^{2}+g^{(4)}_{ij}g^{(4)ij}\Big)+$... (A-16)